OpenStudy (anonymous):

Easy! ;) Please solve, showing all steps? :) 5(x - 3) - 2(2x + 4) = -2(x - 2)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

try the link

OpenStudy (anonymous):

I tried it but it didn't help :/ can u try to do it? :) I would REAALLLYYY appreciate it!! @smartguy1124

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

thanks!! :DDD

OpenStudy (anonymous):

5X-15-4X-8=-2X+4 X-23=-2X+4 3X=27 X=9

OpenStudy (anonymous):

here you are! hope I helped! :) Distribute the 5, -2 amd -2 5x-15-4x-8=-2x+4 Now move the -2x and 4 to the other side by adding/subtracting 5x-15-4x-8+2x-4=-2x+2x+4-4 5x-15-4x-8+2x-4=0 Now collect like terms 3x-27=0 Now add 27 to both sides to get 3x alone 3x-27+27=27 3x=27 Now divide both sides by 3 to get x alone 3x/3=27/3 x=9

OpenStudy (anonymous):

Once you nailed the steps, you'll be good to go, i'll message you later with some tips ;)

OpenStudy (anonymous):

@jishan thanks so much!!!!!!!!!!!!!!! @smartguy1124 i'm going to check that book out! @DaniB WOOOOWW.. LOVE IT!! TO EVERYONE: I wish I could give you all medals!!!!

OpenStudy (anonymous):

you should it really helps

OpenStudy (anonymous):

thanks again!!!!!!!!!! you guys are a really big help.. one of the reasons I still have an A + in math!!! ^_^

OpenStudy (anonymous):

thanks @GabiMB , remember, try try and try again!! It helped me a lot!

OpenStudy (anonymous):

THANKS SISTER.

OpenStudy (anonymous):

thanks bro ;) @jishan

OpenStudy (anonymous):

WELCOME SISTER.